# A quasi-optimal upper bound for induced paths in sparse graphs

### v1

@@ 4 Properties of induced paths
If $b=\ell+1$ then the first assertion trivially holds as there are no vertices of that rank in $G$. Let us thus prove the first assertion assuming $b\leq\ell$, i.e., there is a triangle-vertex of rank $b$ before $u$ in ${\sf triangle}\text{-}{\sf seq}(P)$ with $u$ in its zone. By construction, the root triangle $K$ of this zone ${\sf zone}_{u}(b)$ sends ribs toward each blocking-vertex of rank $b$ in this zone. Moreover, since $b>a$ and so $u\not\in V(K)$, we note that for any vertex $v$ in ${P}{u}$, we have that ${v}{P}{u}$ contains vertices that are not in $K$. Hence ${u}{P}$ may not intersect $K$, as otherwise we would get a chord, which is excluded.
Suppose toward a contradiction that ${u}{P}$ contains a vertex $v$ of rank $b$ in ${\sf zone}_{u}(b)$. Let $w$ be the vertex of $K$ that is contained in $P$. By the above discussion $v$ does not lie in $K$. So $v$ must be the bottom endpoint of a rib of rank $b$. Note that this rib cannot be an edge of $P$ as we would again induce a chord with $w$. Now since the endpoints of $P$ are triangle-vertices, we derive that $v$ must have degree two in $P$. We conclude that $P$ must contain all the representatives of the occurrence of $b$, i.e., the three vertices $x_{1},x_{2},x_{3}$ that were created during the blow-up operation to represent this occurrence of the index $b$ in the barrier; see Section 3.4. However one of these vertices will induce a chord with $w$, a contradiction. This concludes the proof of the first assertion.
Let us now assume toward a contradiction that $P$ contains a triangle-vertex $v$ of rank $i\in\operatorname{{\sf reach}}^{+}_{b}(a)$ contained in ${\sf zone}_{u}(b)$, with possibly $b=\ell+1$ (in which case ${\sf zone}_{u}(b)=V(G)$). As $u$ is correct, every such $i$ is burned, i.e., either $i$ is a ghost rank, or ${P}{u}$ contains the endpoint of a rib of index $i$ incident to some $w\in{\sf zone}\text{-}{\sf ancestor}_{u}(i)$. Clearly, $i$ is not a ghost rank as $v$ belongs to $P$. Thus the second property holds. Note that ${P}{u}$ either contains such a rib of rank $i$, or, using again the fact that the endpoints of $P$ are triangle-vertices, it contains the three representatives of the occurrence of $i$ in the barrier. Both possibilities induced a chord with $v$, a contradiction. This proves the second assertion and hence the proposition. ∎
As already mentioned in the proof of Proposition 4.1, note that after visiting a triangle-vertex $u$, by construction, the path $P$ has several options we review here. Recall that hopping ribs are those additional ribs added between triangles of rank 111 and $j\geq 1$; see Figures 4 and 5. After visiting a triangle-vertex $u$, either:
@@
Moreover, when $u$ is ascending toward a triangle of rank 111 it could be that $P$ contains a hopping rib of the same rank as $u$, whose group of representatives are encountered first in the barrier, in which case we say that $u$ is hopping upmargin: hopping up/down . Similarly, when $u$ is jumping, it could be that the used rib $e$ is a hopping rib, in which case we say that $u$ is hopping down. We distinguish these two cases as they result into $P$ visiting two distinct triangles of a same rank, which, while irrelevant as far as the asymptotic order of a path is concerned, needs extra care in the coming statements.
@@
Let $K$ be the triangle in which $u$ belongs. Note that $u$ may be staying, ascending, descending, or jumping. Let us assume that it is not staying, as otherwise we can replace $u$ by the next vertex in $K$ (at most once) and conduct the same analysis. We will distinguish the remaining cases in the rest of the proof. For each case, we shall show $v\in{\sf zone}_{v}(b)$, and use Proposition 4.7 to show that ${P}{v}$ has visited blocking-vertices of each of the indices required to be burned for $v$ to be correct. Moreover, along the way, we show that ${u}{P}{v}$ is of one of the five types as described in the statement.
@@
Case 1.b: the rank of $K^{\prime}$ is 111; see Figure 8 (right).
@@
Note that since $a,c<b$, $v\in{\sf zone}_{u}(b)$. Also, $c$ is the first occurrence of its index in the trace of ${u}{P}{v}$: this is by definition of the barrier if $c=1$, and in the other case, if $P$ has crossed another occurrence of $c$ before, then one of its attached rib would induce a chord with $v$, a contradiction. So this trace is a prefix of $B(a,1)$ from $a$ to the first occurrence of $c$. Hence it is equal to $B(a,c)$ by definition. By Observation 3.11, this trace contains all indices from $a$ to $c$. Note that $a$ appears as well as the rank of a blocking-vertex in the barrier because of the hopping rib of rank $a$. So each of these indices in burned by $v$. As $a,c<b$, by the transitivity of $\operatorname{{\sf reach}}_{b}$, we have $\operatorname{{\sf reach}}_{b}(c)=\operatorname{{\sf reach}}_{b}(a)$. Moreover every $i\in\operatorname{{\sf reach}}^{+}_{b}(a)$ is burned by $u$. So every $i\in\operatorname{{\sf reach}}^{+}_{b}(c)$ is burned by $v$ and we conclude that $v$ is correct. Note that this situation satisfies the conditions of Case (c).
@@
Case 2.a: ${a=1}$; Figure 9 (left).
Then the trace of the barrier from $u$ to $v$ is the reverse of $B(c,1)$. By Observation 3.11 this barrier contains indices from $c$ to 111, and blocking vertices of rank 222 to $c$, the later being included because of the hopping rib. Note that these indices are greater than 111. Thus no rib can be used by $P$ while crossing this barrier. So $v$ must belong to $K^{\prime}$, and we deduce that $b$ does not appear in the barrier. So $b>c$ and $v$ lies in ${\sf zone}_{u}(b)$. Since $b>c$ and $b>1$, by the transitivity of $\operatorname{{\sf reach}}_{b}$, we have $\operatorname{{\sf reach}}_{b}(1)=\operatorname{{\sf reach}}_{b}(c)$. Now since all $i\in\operatorname{{\sf reach}}^{+}_{b}(1)$ have been burned by $u$, we trivially get that every $i\in\operatorname{{\sf reach}}^{+}_{b}(c)$ is burned by $v$ as well. So $v$ is correct.
@@
The trace of the barrier is ${B(a,a-1)}$. By Observation 3.12, its blocking-vertices are of rank greater than $a$. Hence the indices visited by $P$ belong to $\operatorname{{\sf reach}}^{+}_{b}(a)$. So no rib can be used by $P$ while crossing this barrier, and hence $v$ must belong to $K^{\prime}$. Thus $v\in{\sf zone}_{u}(b)$. Since $v\in{\sf zone}(u)$, $u$ is locking and $a$ is the index-lock of $v$. Let us show that every index in $\operatorname{{\sf reach}}^{+}_{a}(a-1)$ is visited by $P$; we stress the fact that we are now dealing with $a$ as a blocking index, and not $b$. By Proposition 3.15, $a-1$ either lies in ${\sf T}(a^{-})$, or in ${\sf T}(k^{+})$ for some ancestor $k$ of $a$ with possibly $k=a$. Moreover by Proposition 3.8, all indices in $\operatorname{{\sf reach}}_{a}^{+}(a-1)$ are visited, as the trace of ${u}{P}{v}$ contains such an ancestor by Observation 3.5. Hence $v$ is correct.
The trace of ${u}{P}{v}$ is equal to ${B(a,a-1)}$, hence is a factor of $B(\ell)$. Since $u$ is locking the conditions of Case (e) are fulfilled.
@@
Note that $u$ is the top endpoint of a rib having its bottom endpoint in a barrier separating two triangles. Let $K^{s}$ and $K^{t}$ be these triangles corresponding to two adjacent nodes $s,t$ of respective rank $i,j$ with $s$ is the parent of $t$. Note that the trace of the barrier from $s$ to $t$ is either $B(i,j)$ if $i\geq 2$, in which case $j=i-1$, or the reverse of $B(j,1)$, if $i=1$.
@@
The coming arguments hold in these two situations regardless of whether $i=1$ or $i\geq 2$. As the vertices of $K^{s}$ and those in the barrier belong to ${\sf zone}(u)$, we have $c<a$ and $v\in{\sf zone}_{b}(u)$. Since $c<a$, we have $v=\textsf{next}(u)\in{\sf zone}(u)$ and so $u$ is locking. Let us show that every index in $\operatorname{{\sf reach}}^{+}_{a}(c)$ is visited by $P$; note that we are now dealing with $a$ as a blocking index, and not $b$. Since $c<a$, by Proposition 3.15, $c$ either lies in ${\sf T}(a^{-})$, or in ${\sf T}(k^{+})$ for some ancestor $k$ of $a$ with possibly $k=a$. Moreover by Proposition 3.8, all indices in $\operatorname{{\sf reach}}_{a}^{+}(c)$ are visited, as the trace of ${u}{P}{v}$ contains such an ancestor by Observation 3.5. Thus $v$ is correct.
The trace of ${u}{P}{v}$ is a factor of $B(i,j)$, hence of $B(\ell)$. Since $u$ is locking the conditions of Case (e) are fulfilled.
@@
Thus we have $c=j$ in the remaining. Two situations arise depending on whether $i\geq 2$, in which case $j=i-1$, or if $i=1$, in which case $j\geq 1$. If $i\geq 2$, then $v$ also lies in ${\sf zone}_{u}(b)$. The same analysis as in the previous cases yields the desired conclusion with $u$ locking, each of $i\in\operatorname{{\sf reach}}_{a}^{+}(c)$ visited by $P$, $v$ correct, and the trace of ${u}{P}{v}$ being a factor of $B(\ell)$ from $a$ to $c$, i.e., Case (e).
@@
The trace of ${u}{P}{v}$ is the reverse of a prefix of $B(c,1)$ going from the starting occurrence of $c$ to the first occurrence of $a$. Hence this trace is equal to the reverse of $B(c,a)$ by definition. Since $c<a$, $v$ belongs to ${\sf zone}(u)$ and so $u$ is locking. Let us show that every index in $\operatorname{{\sf reach}}^{+}_{a}(c)$ is visited by $P$. Since $c<a$, by Proposition 3.15, $c$ either lies in ${\sf T}(a^{-})$, or in ${\sf T}(k^{+})$ for some ancestor $k$ of $a$ with possibly $k=a$. Moreover by Proposition 3.8, all indices in $\operatorname{{\sf reach}}_{a}^{+}(c)$ are visited, as the trace of ${u}{P}{v}$ contains such an ancestor by Observation 3.5. Thus $v$ is correct.
Note that the trace of ${u}{P}{v}$ is a factor from $c$ to $a$. Hence it satisfies the conditions of Case (e).
@@
The rib incident to $u$ in $P$ must be a hopping rib, as otherwise the hopping ribs of the barrier would induced a chord with $P$. As $u$ is correct and $a=c$, we trivially get that $v$ is correct.
Note that the trace of ${u}{P}{v}$ is $\{a\}$. By construction, $v$ cannot be hopping down, as such ribs are only between consecutive zones of a same rank. Moreover $v$ cannot be hopping up as $P$ would induce a chord with $u$. This corresponds to Case (b).
@@
Then the trace of ${u}{P}{v}$ is the reverse of a prefix of $B(c,1)$ going from the starting occurrence of $c$ to the first occurrence of $a$. Hence this trace is equal to the reverse of $B(c,a)$ by definition. By Observation 3.11 this barrier contains blocking-vertices of ranks from $c-1$ to $a+1$, to which we add $c$ because of the hopping rib. Moreover $b>c$ as otherwise, since $b>a$, we would have $c>b>a$ and by Observation 3.11 the barrier would contain $b$, a contradiction. Since $b>c,a$, by the transitivity of $\operatorname{{\sf reach}}_{b}$, we have $\operatorname{{\sf reach}}_{b}(a)=\operatorname{{\sf reach}}_{b}(c)$. Now since all $i\in\operatorname{{\sf reach}}^{+}_{b}(a)$ have been burned by $u$, clearly every $i\in\operatorname{{\sf reach}}^{+}_{b}(c)$ is burned by $v$. So $v$ is correct.
@@
Note that vertices in ${\sf special}\text{-}{\sf seq}(P)$ have in common that the vertices that follow them in $P$, if they exist, are of a rank smaller than their index-lock: this is trivial for the vertex of largest rank as its index-lock is $\ell+1$, and it follows from Corollary 4.10 for the others. We shall in fact show a stronger statement, using the following notion.
@@
So the trace of $P$ from $u$ to $v$ has total length bounded by $2\cdot|B(k)|$. Note that the repeated indices, in the concatenation, are those of the triangle-vertices visited by $P$. Hence their number is bounded by $p$. Now, since to each index in a barrier corresponds 333 vertices representing it, we derive that $P$ has length at most $3\cdot 2\cdot|B(k)|+p=6\cdot|B(k)|+4k\leq 10\cdot|B(k)|$. ∎
@@
$13\cdot\left(\sum_{i=1}^{p}|B(k_{i})|\right)+1$
@@
We now prove the bound on the order of $P$. Let $v_{1},\dots,v_{p-1}$ be such that $v_{i}:=\textsf{next}(u_{i})$ for all $i\in\left\{1,\dots,p-1\right\}$. Note that $P$ is the concatenation of ${u_{i}}{P}{v_{i}}$ and of ${v_{i}}{P}{u_{i+1}}$ for $i$ ranging from 111 to $p-1$, plus possibly one extra vertex which may be the one sharing the same triangle as $u_{p}$.

### v2

@@ 4 Properties of induced paths
If $b=\ell+1$ then the first assertion trivially holds as there are no vertices of that rank in $G$. Let us thus prove the first assertion assuming $b\leq\ell$, i.e., there is a triangle-vertex of rank $b$ before $u$ in ${\sf triangle}\text{-}{\sf seq}(P)$ with $u$ in its zone. By construction, the triangle $K$ at the root of this zone ${\sf zone}_{u}(b)$ sends ribs toward each blocking-vertex of rank $b$ in this zone. Moreover, since $b>a$ and so $u\not\in V(K)$, we note that for any vertex $v$ in Pu{P}{u}, we have that vPu{v}{P}{u} contains vertices that are not in $K$. Hence uP{u}{P} does not intersect $K$, as otherwise we would get a chord, which is excluded.
Suppose toward a contradiction that uP{u}{P} contains a vertex $v$ of rank $b$ in ${\sf zone}_{u}(b)$. Let $w$ be the vertex of $K$ that is contained in $P$. By the above discussion, $v$ does not lie in $K$. So $v$ must be the bottom endpoint of a rib of rank $b$. Note that this rib cannot be an edge of $P$ as we would again induce a chord with $w$. Now since the endpoints of $P$ are triangle-vertices, we derive that $v$ must have degree two in $P$. We conclude that $P$ must contain all the representatives of the occurrence of $b$, i.e., the three vertices $x_{1},x_{2},x_{3}$ that were created during the blow-up operation to represent this occurrence of the index $b$ in the barrier; see Section 3.4. However one of these vertices will induce a chord with $w$, a contradiction. This concludes the proof of the first assertion.
Let us now assume toward a contradiction that $P$ contains a triangle-vertex $v$ of rank $i\in\operatorname{{\sf reach}}^{+}_{b}(a)$ contained in ${\sf zone}_{u}(b)$, with possibly $b=\ell+1$ (in which case ${\sf zone}_{u}(b)=V(G)$). As $u$ is correct, every such $i$ is burned, i.e., either $i$ is a ghost rank, or Pu{P}{u} contains the endpoint of a rib of index $i$ incident to some $w\in{\sf zone}\text{-}{\sf ancestor}_{u}(i)$. Clearly, $i$ is not a ghost rank as $v$ belongs to $P$. Thus the second property holds. Note that Pu{P}{u} either contains such a rib of rank $i$, or, using again the fact that the endpoints of $P$ are triangle-vertices, it contains the three representatives of the occurrence of $i$ in the barrier. Both possibilities induced a chord with $v$, a contradiction. This proves the second assertion and hence the proposition. ∎
As already mentioned in the proof of Proposition 4.1, note that after visiting a triangle-vertex $u$, by construction, the path $P$ has several options we review here. Recall that hopping ribs are those additional ribs added between triangles of rank $1$ and $j\geq 1$; see Figures 4 and 5. After visiting a triangle-vertex $u$, either:
@@
Moreover, when $u$ is ascending toward a triangle of rank $1$ it could be that $P$ contains a hopping rib of the same rank as $u$, whose group of representatives are encountered first in the barrier, in which case we say that $u$ is hopping upmargin: hopping up/down . Similarly, when $u$ is jumping, it could be that the used rib $e$ is a hopping rib, in which case we say that $u$ is hopping down. We distinguish these two cases as they result into $P$ visiting two distinct triangles of a same rank, which, while irrelevant as far as the asymptotic order of a path is concerned, needs extra care in the coming statements.
@@
Let $K$ be the triangle in which $u$ belongs. Note that $u$ may be staying, ascending, descending, or jumping. Let us assume that it is not staying, as otherwise we can replace $u$ by the next vertex in $K$ (at most once) and conduct the same analysis. We will distinguish the remaining cases in the rest of the proof. For each case, we shall show $v\in{\sf zone}_{v}(b)$, and use Proposition 4.7 to show that Pv{P}{v} has visited blocking-vertices of each of the indices required to be burned for $v$ to be correct. Moreover, along the way, we show that uPv{u}{P}{v} is of one of the five types as described in the statement.
@@
Case 1.b: the rank of $K^{\prime}$ is $1$; see Figure 8 (right).
@@
Note that since $a,c<b$, $v\in{\sf zone}_{u}(b)$. Also, $c$ is the first occurrence of its index in the trace of uPv{u}{P}{v}: this is by definition of the barrier if $c=1$, and in the other case, if $P$ has crossed another occurrence of $c$ before, then one of its attached rib would induce a chord with $v$, a contradiction. So this trace is a prefix of $B(a,1)$ from $a$ to the first occurrence of $c$. Hence it is equal to $B(a,c)$ by definition. By Observation 3.11, this trace contains all indices from $a$ to $c$. Note that $a$ appears as well as the rank of a blocking-vertex in the barrier because of the hopping rib of rank $a$. So each of these indices in burned by $v$. As $a,c<b$, by the transitivity of $\operatorname{{\sf reach}}_{b}$, we have $\operatorname{{\sf reach}}_{b}(c)=\operatorname{{\sf reach}}_{b}(a)$. Moreover every $i\in\operatorname{{\sf reach}}^{+}_{b}(a)$ is burned by $u$. So every $i\in\operatorname{{\sf reach}}^{+}_{b}(c)$ is burned by $v$ and we conclude that $v$ is correct. Note that this situation satisfies the conditions of Case (c).
@@
Case 2.a: a=1{a=1}; Figure 9 (left).
Then the trace of the barrier from $u$ to $v$ is the reverse of $B(c,1)$. By Observation 3.11 this barrier contains indices from $c$ to $1$, and blocking vertices of rank $2$ to $c$, the later being included because of the hopping rib. Note that these indices are greater than $1$. Thus no rib can be used by $P$ while crossing this barrier. So $v$ must belong to $K^{\prime}$, and we deduce that $b$ does not appear in the barrier. So $b>c$ and $v$ lies in ${\sf zone}_{u}(b)$. Since $b>c$ and $b>1$, by the transitivity of $\operatorname{{\sf reach}}_{b}$, we have $\operatorname{{\sf reach}}_{b}(1)=\operatorname{{\sf reach}}_{b}(c)$. Now since all $i\in\operatorname{{\sf reach}}^{+}_{b}(1)$ have been burned by $u$, we trivially get that every $i\in\operatorname{{\sf reach}}^{+}_{b}(c)$ is burned by $v$ as well. So $v$ is correct.
@@
The trace of the barrier is B(a,a1){B(a,a-1)}. By Observation 3.12, its blocking-vertices are of rank greater than $a$. Hence the indices visited by $P$ belong to $\operatorname{{\sf reach}}^{+}_{b}(a)$. So no rib can be used by $P$ while crossing this barrier, and hence $v$ must belong to $K^{\prime}$. Thus $v\in{\sf zone}_{u}(b)$. Since $v\in{\sf zone}(u)$, $u$ is locking and $a$ is the index-lock of $v$.
Let us show that every index in $\operatorname{{\sf reach}}^{+}_{a}(a-1)$ is visited by $P$; we stress the fact that we are now dealing with $a$ as a blocking index, and not $b$. By Proposition 3.15, a1a-1 either lies in ${\sf T}(a^{-})$, or in ${\sf T}(k^{+})$ for some ancestor $k$ of $a$ with possibly $k=a$. Moreover by Proposition 3.8, all indices in $\operatorname{{\sf reach}}_{a}^{+}(a-1)$ are visited, as the trace of uPv{u}{P}{v} contains such an ancestor by Observation 3.5. Hence $v$ is correct.
The trace of uPv{u}{P}{v} is equal to B(a,a1){B(a,a-1)}, hence is a factor of $B(\ell)$. Since $u$ is locking the conditions of Case (e) are fulfilled.
@@
Note that $u$ is the top endpoint of a rib having its bottom endpoint in a barrier separating two triangles. Let $K^{s}$ and $K^{t}$ be these triangles corresponding to two adjacent nodes $s,t$ of respective rank $i,j$ with $s$ is the parent of $t$. Note that the trace of the barrier from $s$ to $t$ is either $B(i,j)$ if $i\geq 2$, in which case j=i1j=i-1, or the reverse of $B(j,1)$, if $i=1$.
@@
The coming arguments hold in these two situations regardless of whether $i=1$ or $i\geq 2$. As the vertices of $K^{s}$ and those in the barrier belong to ${\sf zone}(u)$, we have $c<a$ and $v\in{\sf zone}_{b}(u)$. Since $c<a$, we have $v=\textsf{next}(u)\in{\sf zone}(u)$ and so $u$ is locking. Let us show that every index in $\operatorname{{\sf reach}}^{+}_{a}(c)$ is visited by $P$; note that we are now dealing with $a$ as a blocking index, and not $b$. Since $c<a$, by Proposition 3.15, $c$ either lies in ${\sf T}(a^{-})$, or in ${\sf T}(k^{+})$ for some ancestor $k$ of $a$ with possibly $k=a$. Moreover by Proposition 3.8, all indices in $\operatorname{{\sf reach}}_{a}^{+}(c)$ are visited, as the trace of uPv{u}{P}{v} contains such an ancestor by Observation 3.5. Thus $v$ is correct.
The trace of uPv{u}{P}{v} is a factor of $B(i,j)$, hence of $B(\ell)$. Since $u$ is locking the conditions of Case (e) are fulfilled.
@@
Thus we have $c=j$ in the remaining. Two situations arise depending on whether $i\geq 2$, in which case j=i1j=i-1, or if $i=1$, in which case $j\geq 1$. If $i\geq 2$, then $v$ also lies in ${\sf zone}_{u}(b)$. The same analysis as in the previous cases yields the desired conclusion with $u$ locking, each of $i\in\operatorname{{\sf reach}}_{a}^{+}(c)$ visited by $P$, $v$ correct, and the trace of uPv{u}{P}{v} being a factor of $B(\ell)$ from $a$ to $c$, i.e., Case (e).
@@
The trace of uPv{u}{P}{v} is the reverse of a prefix of $B(c,1)$ going from the starting occurrence of $c$ to the first occurrence of $a$. Hence this trace is equal to the reverse of $B(c,a)$ by definition. Since $c<a$, $v$ belongs to ${\sf zone}(u)$ and so $u$ is locking. Let us show that every index in $\operatorname{{\sf reach}}^{+}_{a}(c)$ is visited by $P$. Since $c<a$, by Proposition 3.15, $c$ either lies in ${\sf T}(a^{-})$, or in ${\sf T}(k^{+})$ for some ancestor $k$ of $a$ with possibly $k=a$. Moreover by Proposition 3.8, all indices in $\operatorname{{\sf reach}}_{a}^{+}(c)$ are visited, as the trace of uPv{u}{P}{v} contains such an ancestor by Observation 3.5. Thus $v$ is correct.
Note that the trace of uPv{u}{P}{v} is a factor from $c$ to $a$. Hence it satisfies the conditions of Case (e).
@@
The rib incident to $u$ in $P$ must be a hopping rib, as otherwise the hopping ribs of the barrier would induce a chord with $P$. As $u$ is correct and $a=c$, we trivially get that $v$ is correct.
Note that the trace of uPv{u}{P}{v} is $\{a\}$. By construction, $v$ cannot be hopping down, as such ribs are only between consecutive zones of a same rank. Moreover $v$ cannot be hopping up as $P$ would induce a chord with $u$. This corresponds to Case (b).
@@
Then the trace of uPv{u}{P}{v} is the reverse of a prefix of $B(c,1)$ going from the starting occurrence of $c$ to the first occurrence of $a$. Hence this trace is equal to the reverse of $B(c,a)$ by definition. By Observation 3.11 this barrier contains blocking-vertices of ranks from c1c-1 to $a+1$, to which we add $c$ because of the hopping rib. Moreover $b>c$ as otherwise, since $b>a$, we would have $c>b>a$ and by Observation 3.11 the barrier would contain $b$, a contradiction. Since $b>c,a$, by the transitivity of $\operatorname{{\sf reach}}_{b}$, we have $\operatorname{{\sf reach}}_{b}(a)=\operatorname{{\sf reach}}_{b}(c)$. Now since all $i\in\operatorname{{\sf reach}}^{+}_{b}(a)$ have been burned by $u$, clearly every $i\in\operatorname{{\sf reach}}^{+}_{b}(c)$ is burned by $v$. So $v$ is correct.
@@
Note that vertices in ${\sf special}\text{-}{\sf seq}(P)$ have in common that the vertices that follow them in $P$, if they exist, are of a rank smaller than their index-lock: this is trivial for the vertex of largest rank as its index-lock is $\ell+1$, and it follows from Corollary 4.10 for the others. We shall in fact show a stronger statement, using the following notion. Recall that ${\sf T}(i)$ denotes the subtree of $\sf T$} rooted at index $i$.
@@
So the trace of $P$ from $u$ to $v$ has total length bounded by $2\cdot|B(k)|$. Note that the repeated indices, in the concatenation, are those of the triangle-vertices visited by $P$. Hence their number is bounded by $p$. Now, since to each index in a barrier corresponds $3$ vertices representing it, we derive that $P$ has length at most $3\cdot 2\cdot|B(k)|+p=6\cdot|B(k)|+4k\leq 10\cdot|B(k)|$. ∎
@@
$13\cdot\left(\sum_{i=1}^{p}|B(k_{i})|\right)+1.$
@@
We now prove the bound on the order of $P$. Let $v_{1},\dots,v_{p-1}$ be such that $v_{i}:=\textsf{next}(u_{i})$ for all $i\in\left\{1,\dots,p-1\right\}$. Note that $P$ is the concatenation of ${u_{i}}{P}{v_{i}}$ and of ${v_{i}}{P}{u_{i+1}}$ for $i$ ranging from $1$ to p1p-1, plus possibly one extra vertex which may be the one sharing the same triangle as $u_{p}$.